OpenStudy (anonymous):

A satellite , B, is 17,000 miles from the horizon of earth's radius is about 4,000 miles. Find the approximate distance the satellite is from the earth's surface. -13,464 miles -17,464 miles -20,957 miles -21,000 miles

OpenStudy (anonymous):

Help Help Help <3

OpenStudy (nurali):

Note: all numbers are in thousands (so when I type 4, I mean "4 thousand") AO = 4 AE = 17 EO = x+4 --------------- AE^2 + AO^2 = EO^2 .... pythagorean theorem 17^2 + 4^2 = (x+4)^2 289+16 = (x+4)^2 305 = (x+4)^2 (x+4)^2 = 305 x+4 = sqrt(305) x = -4+sqrt(305) x = -4+17.464 x = 13.464 Remember I stated that all values are in thousands, so x = 13.464 really means 13,464 So the approximate distance the satellite is from the Earth’s surface is 13,464 miles

OpenStudy (anonymous):

Thank you! :D

OpenStudy (nurali):

Anytime.